OpenStudy (anonymous):

Find two functions defined implicitly by the given relation: x^2+y^2=36

OpenStudy (aum):

\[ x^2+y^2=36 \\ y^2 = 36 - x^2 \\ y = \pm \sqrt{36-x^2} \\ y = +\sqrt{36-x^2} ~~\text{ and }~~y = -\sqrt{36-x^2} \]

OpenStudy (confluxepic):

Correct @aum